Required information

[The following information applies to the questions displayed below.]

Saddleback Company makes camping lanterns using a single production process. All direct materials are added at the beginning of the manufacturing process. Information for the month of March follows:

Units Costs
Beginning work in process (30% complete) 127,800
Direct materials $ 182,000
Conversion cost 334,000
Total cost of beginning work in process $ 516,000
Number of units started 233,000
Number of units completed and transferred to finished goods 333,800
Ending work in process (65% complete) ?
Current period costs
Direct materials $ 497,400
Conversion cost 638,000
Total current period costs $ 1,135,400

Required:

1. Using the weighted-average method of process costing, complete each of the following steps:

a. Reconcile the number of physical units worked on during the period.

Physical Units Physical Units
Beginning Units Units Completed
Units Started Ending Units
Total Units 0 Total Units 0

b. Calculate the number of equivalent units.

Equivalent Units
Physical Units Direct Materials Conversion
Units Completed
Ending Inventory
Total 0 0 0

c. Calculate the cost per equivalent unit.

Direct Materials Conversion
Cost per Equivalent Unit

d. Reconcile the total cost of work in process.

Direct Materials Conversion Total Cost
Units Completed
Ending Inventory
Total Cost Accounted for $0 $0 $0
Answer #1
1a
Physical Units Physical Units
Beginning Units 127800 Units Completed 333800
Units Started 233000 Ending Units 27000
Total Units 360800 Total Units 360800
1b
Equivalent Units
Physical Units Direct Materials Conversion
Units Completed 333800 333800 333800
Ending Inventory 27000 27000 17550
Total 360800 360800 351350
c
Direct Materials Conversion
Beginning work in process costs 182000 334000
Current period costs 497400 638000
Total costs 679400 972000
Divide by Equivalent units 360800 351350
Cost per Equivalent Unit 1.88304 2.76647
d
Direct Materials Conversion Total cost
Units Completed 628559 923448 1552007
Ending Inventory 50842 48552 99394
Total Cost Accounted for 679401 972000 1651401
Workings for d:
Direct Materials Conversion
Units Completed =333800*1.88304 =333800*2.76647
Ending Inventory =27000*1.88304 =17550*2.76647